OpenStudy (anonymous):

The ratio of the lengths of corresponding sides of two similar triangles is 7:10. The smaller triangle has an area of 63 cm2. What is the area of the larger triangle? 70 cm2 630 cm2 90 cm2 128.6 cm2

OpenStudy (crashonce):

Alright to do this question you need to know this simple formula: \[\frac{ l1^2 }{ l2^2 } = \frac{ A1 }{ A2 }\] where l1 and l2 are the sides of the triangles and A1 and A2 are the areas

OpenStudy (anonymous):

Okay so it would be 90??

OpenStudy (crashonce):

so that means that \[\frac{ 7^2 }{ 10^2 } = \frac{ 63 }{ A2 }\\]

OpenStudy (crashonce):

no

OpenStudy (crashonce):

using the formula i just typed in you can see the LHS will be 49/100 and the RHS will be 63/A2 Therefore, simplifying, we get 49A2 = 6300 THerefore A2= 128.6 D is the answer

OpenStudy (anonymous):

Oh thanks, sorry I had to go eat dinner.

OpenStudy (crashonce):

np